### v1

# Photon Statistics for Fock and Coherent States Interfering in a Beamsplitter

Abstract
[…]
Keywords
quantum optics, photon statistics, beamsplitter, q parameter, second-order coherence function.
1 Introduction
Quantum optics is a field dedicated to studying the quantum nature of light and its interaction with matter [1, 2]. It plays a fundamental role in the development of emerging quantum technologies such as quantum communication [3], quantum computation [4, 5], high-precision metrology [6], and quantum sensing [7]. The manipulation and control of quantum states of light are essential for these applications, where phenomena like quantum interference, entanglement, and squeezing are routinely exploited [8, 9].
The beamsplitter (BS) is one of the most basic yet powerful devices employed in quantum optics laboratories and experiments [10]. It enables the coherent superposition of optical modes, leading to interference effects that are central to protocols such as the Hong–Ou–Mandel effect [11] and quantum state tomography [12]. Understanding how various quantum states of light transform through a BS is therefore critical, both for theoretical modeling and experimental implementations.
Over the past decades, numerous experiments have explored the behavior of single-photons, multi-photon Fock states, coherent states, and their superpositions passing through BSs [13, 14]. These studies have revealed rich photon statistics, nonclassical correlations, and phase space interference patterns. Notably, the observation of negative regions in the Wigner function serves as a hallmark of nonclassicality and has been confirmed in experiments involving Fock states and Schrödinger cat states [15, 16].
A key criterion to characterize the photon number statistics of a quantum state of light is Mandel’s $Q$ parameter [17], defined by the normalized variance of the photon number distribution relative to a Poisson distribution. This parameter quantifies deviations from Poissonian statistics: $Q=0$ corresponds to a Poissonian distribution (typical of coherent states), $Q>0$ indicates super-Poissonian (bunched) statistics typical of thermal light, and $Q<0$ reveals sub-Poissonian statistics, a signature of nonclassical light exhibiting photon antibunching [22]. Mandel’s $Q$ parameter thus provides a straightforward guideline to assess the quantum nature of light sources.
Closely related to $Q$ is the second-order coherence function $g^{(2)}(\tau)$, introduced by Glauber [1], which measures temporal correlations between photon detection events. At zero time delay, $g^{(2)}(0)$ can be expressed in terms of the photon number variance and mean, and thus relates directly to Mandel’s $Q$ parameter. This relation allows one to interpret the photon bunching or antibunching behavior of the field in terms of its number statistics, as we will see later.
[…]
This work is not intended to provide an exhaustive or highly technical analysis of all possible input-output configurations of BSs. Instead, we focus on fundamental states that allow a clear illustration of photon statistics and the role of quantum versus classical states of light. In this sense, the present discussion is meant to serve as a pedagogical guide, complementing standard quantum optics references and providing an accessible starting point for students and researchers interested in exploring more advanced topics, including quantum correlations and coherence theory.
1.1 Beamsplitter Theory
A bi-modal BS is a fundamental linear optical apparatus that mixes two input modes of the electromagnetic field into two output modes through a unitary transformation [23, 24]. In the quantum description, the BS acts on the bosonic creation operators $a_{a}^{\dagger}$ and $a_{b}^{\dagger}$ associated with the input modes $a$ and $b$, transforming them into output mode operators $a_{c}^{\dagger}$ and $a_{d}^{\dagger}$, as seen in Fig. (1).
Figure 1: Scheme of a lossless beamsplitter.
A common and widely used model for a lossless BS with complex amplitude reflection coefficient $r$ and transmission coefficient $t$, satisfying [23]
$|r|=|r^{\prime}|,\quad|t|=|t^{\prime}|,\quad|r|^{2}+|t|^{2}=1,$
$r^{*}t+r^{\prime}t^{*}=0,\quad r^{*}t+r^{\prime}t^{\prime*}=1,$
is given by the transformation:
$a_{c}=ta_{a}+ira_{b},\quad a_{d}=ira_{a}+ta_{b},$ (1.1)
where the factor $i$ in the reflection terms ensures the overall unitary operation and proper phase relations [25].
The inverse transformation reads:
$a_{a}=t^{*}a_{c}-ir^{*}a_{d},\quad a_{b}=-ir^{*}a_{c}+t^{*}a_{d},$ (1.2)
which preserves the bosonic commutation relations:
$a_{i},a_{j}^{\dagger}]=\delta_{ij},\quad i,j=a,b,c,d.$
The unitary operator $U_{BS}$ acts on the two-mode space, such that
$a_{c}=U_{BS}^{\dagger}a_{a}U_{BS},\quad a_{d}=U_{BS}^{\dagger}a_{b}U_{BS},$
where
$U_{BS}(\theta,\phi)=\exp\left[\theta\left(e^{i\phi}a_{a}^{\dagger}a_{b}-e^{-i\phi}a_{a}a_{b}^{\dagger}\right)\right],$ (1.3)
with
  • $\theta\in[0,\pi/2]$ defines the mixing angle, with the amplitude transmission and reflection coefficients given by $|t|=\cos\theta$ and $|r|=\sin\theta$,
  • $\phi$ is a relative phase shift introduced between the reflected and transmitted fields.
1.2 Basic Photon Statistics
The photon number operators for the output modes are:
$\mathcal{N}_{k}=a_{k}^{\dagger}a_{k},\quad k=c,d.$ (1.4)
For a general state $\rho$, the mean photon number at each output is [26]:
$\langle\mathcal{N}_{k}\rangle=\mathrm{Tr}\left(\rho\mathcal{N}_{k}\right).$ (1.5)
The photon number variance is:
$(\Delta\mathcal{N}_{k})^{2}=\langle\mathcal{N}_{k}^{2}\rangle-\langle\mathcal{N}_{k}\rangle^{2},$ (1.6)
where $\langle\mathcal{N}_{k}^{2}\rangle=\mathrm{Tr}\left(\rho\mathcal{N}_{k}^{2}\right).$
Mandel’s $Q$ parameter, which measures the deviation from Poissonian statistics, is defined as [26]:
$Q_{k}=\frac{(\Delta\mathcal{N}_{k})^{2}-\langle\mathcal{N}_{k}\rangle}{\langle\mathcal{N}_{k}\rangle}.$ (1.7)
In practice, these quantities are computed by expressing $\mathcal{N}_{k}$ and $\mathcal{N}_{k}^{2}$ in terms of the input operators $a_{a},a_{b}$ using the BS transformations, and evaluating their expectation values with the chosen input state.
Another important quantity is the second-order coherence function at zero delay, $g_{k}^{(2)}(0)$, which characterizes the photon correlation statistics. It is defined as:
$g^{(2)}_{k}(0)==\frac{\langle\mathcal{N}_{k}^{2}\rangle-\langle\mathcal{N}_{k}\rangle}{\langle\mathcal{N}_{k}\rangle^{2}}.$ (1.8)
The $g^{(2)}(0)$ parameter quantifies the conditional probability of detecting two photons simultaneously at the same output mode, normalized to the square of the mean photon number. It provides further insight into the quantum or classical nature of the light field: $g^{(2)}(0)<1$ indicates antibunching (nonclassical light), $g^{(2)}(0)=1$ corresponds to Poissonian (coherent) statistics, and $g^{(2)}(0)>1$ reveals bunching (typically thermal or chaotic light) [1].
Moreover, there is a direct relationship between the Mandel $Q$ parameter and $g^{(2)}(0)$:
$g^{(2)}_{k}(0)=1+\frac{Q_{k}}{\langle\mathcal{N}_{k}\rangle}.$ (1.9)
2 Fock State Inputs $|n\rangle|m\rangle$
A detailed treatment of Fock-state inputs in a beamsplitter can be found in the seminal work of Campos, Saleh, and Teich [25] and chapter 5 of Agarwal’s textbook [22]. Following their approach, we consider a two-mode input Fock state of the type
$|in\rangle=|n\rangle_{a}|m\rangle_{b}=\frac{(a_{a}^{\dagger})^{n}}{\sqrt{n!}}\frac{(a_{b}^{\dagger})^{m}}{\sqrt{m!}}|0\rangle_{a}|0\rangle_{b},$ (2.1)
where $|n\rangle_{a}$ and $|m\rangle_{b}$ are photon-number eigenstates in the input ports.
The output state after passing through the BS is
$|out\rangle$ $=U_{BS}|n\rangle_{a}|m\rangle_{b}$
$=\frac{(ta_{c}^{\dagger}+ira_{d}^{\dagger})^{n}}{\sqrt{n!}}\frac{(ira_{c}^{\dagger}+ta_{d}^{\dagger})^{m}}{\sqrt{m!}}|0\rangle_{c}|0\rangle_{d}.$ (2.2)
To obtain the explicit form of the output state in the photon-number basis, we expand each factor using the multi-binomial theorem:
$(ta_{c}^{\dagger}+ira_{d}^{\dagger})^{n}$ $=\sum_{p=0}^{n}\binom{n}{p}t^{n-p}(ir)^{p}(a_{c}^{\dagger})^{n-p}(a_{d}^{\dagger})^{p},$ (2.3)
$(ira_{c}^{\dagger}+ta_{d}^{\dagger})^{m}$ $=\sum_{q=0}^{m}\binom{m}{q}(ir)^{m-q}t^{q}(a_{c}^{\dagger})^{m-q}(a_{d}^{\dagger})^{q}.$ (2.4)
Multiplying these expansions, we find
$|out\rangle$ $=\frac{1}{\sqrt{n!m!}}\sum_{p=0}^{n}\sum_{q=0}^{m}\binom{n}{p}\binom{m}{q}t^{n-p+q}(ir)^{m-q+p}$
$\quad\times(a_{c}^{\dagger})^{(n-p)+(m-q)}(a_{d}^{\dagger})^{p+q}|0\rangle_{c}|0\rangle_{d}.$ (2.5)
Since
$(a_{c}^{\dagger})^{k}(a_{d}^{\dagger})^{l}|0\rangle_{c}|0\rangle_{d}=\sqrt{k!l!}|k\rangle_{c}|l\rangle_{d},$ (2.6)
the final expression for the output state is
$|out\rangle$ $=\sum_{p=0}^{n}\sum_{q=0}^{m}\frac{\binom{n}{p}\binom{m}{q}}{\sqrt{n!m!}}t^{n-p+q}(ir)^{m-q+p}$
$\quad\times\sqrt{(n-p+m-q)!(p+q)!}$
$\quad\times|n-p+m-q\rangle_{c}|p+q\rangle_{d}.$ (2.7)
Equation (2.7) shows that the output is a superposition of two-mode Fock states with all possible photon-number partitions between the two output ports.
2.1 Mean Photon Number
Using the BS transformations from equation (1.1), we can express the photon number operators at the output as:
$\mathcal{N}_{c}=a_{c}^{\dagger}a_{c}$ $=|t|^{2}a_{a}^{\dagger}a_{a}+|r|^{2}a_{b}^{\dagger}a_{b}+itr^{*}a_{a}^{\dagger}a_{b}-it^{*}ra_{b}^{\dagger}a_{a}$
$=|t|^{2}\mathcal{N}_{a}+|r|^{2}\mathcal{N}_{b}+itr^{*}a_{a}^{\dagger}a_{b}-it^{*}ra_{b}^{\dagger}a_{a},$ (2.8)
$\mathcal{N}_{d}=a_{d}^{\dagger}a_{d}$ $=|r|^{2}a_{a}^{\dagger}a_{a}+|t|^{2}a_{b}^{\dagger}a_{b}-itr^{*}a_{a}^{\dagger}a_{b}+it^{*}ra_{b}^{\dagger}a_{a}$
$=|r|^{2}\mathcal{N}_{a}+|t|^{2}\mathcal{N}_{b}-itr^{*}a_{a}^{\dagger}a_{b}+it^{*}ra_{b}^{\dagger}a_{a}.$ (2.9)
Since the input state is diagonal in the Fock basis, cross terms such as $\langle a_{a}^{\dagger}a_{b}\rangle$ vanish. And, as we know:
$\mathcal{N}_{\gamma}|\gamma\rangle=\gamma|\gamma\rangle,$
hence, the mean photon numbers are the well-known relations:
$\langle\mathcal{N}_{c}\rangle=n|t|^{2}+m|r|^{2},\quad\langle\mathcal{N}_{d}\rangle=n|r|^{2}+m|t|^{2}.$ (2.10)
Note that the total photon number is conserved:
$\langle\mathcal{N}_{c}\rangle+\langle\mathcal{N}_{d}\rangle=n+m.$ (2.11)
2.2 Photon Number Variance
To compute the variance $(\Delta\mathcal{N}_{c})^{2}$, we expand $\mathcal{N}_{c}^{2}=(a_{c}^{\dagger}a_{c})^{2}$, with:
$\langle a_{a}^{\dagger}a_{a}\rangle=n,\quad\langle(a_{a}^{\dagger}a_{a})^{2}\rangle=n^{2},$ (2.12)
$\langle a_{b}^{\dagger}a_{b}\rangle=m,\quad\langle(a_{b}^{\dagger}a_{b})^{2}\rangle=m^{2},$ (2.13)
therefore, from equation (2.8):
$\langle\mathcal{N}^{2}_{c}\rangle$ $=\Big{\langle}\Big{(}|t|^{2}n+|r|^{2}m+itr^{*}a^{\dagger}_{a}a_{b}-it^{*}ra^{\dagger}_{b}a_{a}\Big{)}^{2}\Big{\rangle}$
$=|t|^{4}n^{2}+|r|^{4}m^{2}+2|t|^{2}|r|^{2}nm$
$\quad+\Big{\langle}|t|^{2}|r|^{2}\Big{(}a^{\dagger}_{a}\underbrace{a_{b}a^{\dagger}_{b}}_{\mathcal{N}_{b}+1}a_{a}+a^{\dagger}_{b}\underbrace{a_{a}a^{\dagger}_{a}}_{\mathcal{N}_{a}+1}a_{b}\Big{)}\Big{\rangle}$
$=\langle\mathcal{N}_{c}\rangle^{2}+|t|^{2}|r|^{2}\Big{(}(m+1)n+(n+1)m\Big{)}$
$=\langle\mathcal{N}_{c}\rangle^{2}+|t|^{2}|r|^{2}(2nm+n+m).$ (2.14)
The variance is then [25]:
$(\Delta\mathcal{N}_{c})^{2}=|t|^{2}|r|^{2}(n+m+2nm).$ (2.15)
2.3 Mandel’s $Q$ Parameter
The Mandel $Q$ parameter for the $c$-mode is:
$Q_{c}=\frac{(\Delta\mathcal{N}_{c})^{2}-\langle\mathcal{N}_{c}\rangle}{\langle\mathcal{N}_{c}\rangle}.$ (2.16)
Substituting the expressions above:
$Q_{c}=\frac{|t|^{2}|r|^{2}(n+m+2nm)}{n|t|^{2}+m|r|^{2}}-1,\ \text{for }n=m\neq 0,$ (2.17)
since vacuum is coherent, $Q_{c}=0$.
2.4 Second-order coherence function $g^{2}(0)$
With equations (2.17) and (2.10) in (1.9), we have for mode $c$:
$g^{2}_{c}(0)=1+\frac{1}{n|t|^{2}+m|r|^{2}}\Bigg{(}\frac{|t|^{2}|r|^{2}(n+m+2nm)}{n|t|^{2}+m|r|^{2}}-1\Bigg{)}.$ (2.18)
Both parameters from equations (2.17) and (2.18) are plotted in figures (2) and (3), respectively.
Figure 2: Mandel $Q_{c}$ parameter for mode $c$ of a 50:50 BS with Fock-Fock state input of the type $|n\rangle_{a}|m\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$. (a) Negative $Q_{c}$ values certify nonclassicality [22], demonstrating sub-Poissonian statistics from the input Fock states’ zero photon-number variance $(\Delta\mathcal{N})^{2}=0$. (b) At $m=n=1$, $Q_{c}=0$ despite the output being quantum (e.g., a HOM state), illustrating that $Q\geq 0$ does not guarantee classicality. (c) As $m\gg n$, $Q_{c}$ increases due to dominant super-Poissonian statistics. Numerical parameters $n\in\{0,1,2,5,10\}$ shows how input asymmetry affects quantum features.
Figure 3: $g^{(2)}(0)$ for mode $c$ of a 50:50 BS with Fock-Fock input of the type $|n\rangle_{a}|m\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$. Notably, the Hong-Ou-Mandel (HOM) effect is observed for $n=m=1$, where perfect quantum interference between two indistinguishable photons leads to $g^{(2)}(0)=0$, indicating the complete suppression of simultaneous photon detections at the output.
2.5 Special Cases
2.5.1 Single-photon input $|1\rangle_{a}|0\rangle_{b}$.
When $n=1$ and $m=0$, equation (2.7) reduces to:
$|out\rangle_{cd}=t|1\rangle_{c}|0\rangle_{d}+ir|0\rangle_{c}|1\rangle_{d},$ (2.19)
which describes a single-photon in a superposition between the two output modes. Its statistics is:
$\langle\mathcal{N}_{c}\rangle=|t|^{2},\ (\Delta\mathcal{N}_{c})^{2}=|t|^{2}|r|^{2},\ Q_{c}=-|t|^{2},\ g^{2}_{c}(0)=0.$
2.5.2 One photon in each input $|1\rangle_{a}|1\rangle_{b}$
For $n=m=1$,
$|out\rangle_{cd}=irt\sqrt{2}|2,0\rangle_{cd}+(t^{2}-r^{2})|1,1\rangle_{cd}+irt\sqrt{2}|0,2\rangle_{cd}.$ (2.20)
In the special case of a 50:50 BS ($|t|=|r|=1/\sqrt{2}$), the term $|1\rangle_{c}|1\rangle_{d}$ vanishes due to destructive interference, leading to the well-known Hong-Ou-Mandel state:
$|out\rangle_{cd}=\frac{i}{\sqrt{2}}\Big{(}|2,0\rangle_{cd}+|0,2\rangle_{cd}\Big{)},$ (2.21)
up to a phase. With photon statistics being:
$\langle\mathcal{N}_{c}\rangle=\langle\mathcal{N}_{d}\rangle$ $=1,\quad(\Delta\mathcal{N}_{c})^{2}=(\Delta\mathcal{N}_{d})^{2}=1,$
$\quad\ Q_{c}=Q_{d}$ $=0,\quad g^{2}_{c}(0)=g^{2}_{d}(0)=1.$
These results are well-known in quantum optics [23].
3 Hybrid Inputs $|n\rangle_{a}|\alpha\rangle_{b}$
We now consider a hybrid state where mode $a$ is a Fock state $|n\rangle$ and mode $b$ a coherent state $|\alpha\rangle$.
$|in\rangle_{ab}=|n\rangle_{a}|\alpha\rangle_{b}=\frac{(a_{a}^{\dagger})^{n}}{\sqrt{n!}}|0\rangle_{a}\otimes\mathcal{D}_{b}(\alpha)|0\rangle_{b},$ (3.1)
where $\mathcal{D}_{b}(\alpha)=\exp(\alpha a_{b}^{\dagger}-\alpha^{*}a_{b})$ is the displacement operator for mode $b$.
The output state after the BS is:
$|out\rangle_{cd}$ $=U_{BS}\frac{(a_{a}^{\dagger})^{n}}{\sqrt{n!}}|0\rangle_{a}|\alpha\rangle_{b}$
$=\frac{(ta_{c}^{\dagger}+ira_{d}^{\dagger})^{n}}{\sqrt{n!}}\mathcal{D}_{c}(ir\alpha)\mathcal{D}_{d}(t\alpha)|0\rangle_{c}|0\rangle_{d}.$ (3.2)
Expanding the polynomial:
$(ta_{c}^{\dagger}+ira_{d}^{\dagger})^{n}=\sum_{k=0}^{n}\binom{n}{k}t^{n-k}(ir)^{k}(a_{c}^{\dagger})^{n-k}(a_{d}^{\dagger})^{k}.$ (3.3)
The output state is:
$|out\rangle_{cd}$ $=\sum_{k=0}^{n}\binom{n}{k}\frac{t^{n-k}(ir)^{k}}{\sqrt{n!}}(a_{c}^{\dagger})^{n-k}(a_{d}^{\dagger})^{k}|ir\alpha\rangle_{c}|t\alpha\rangle_{d}.$ (3.4)
3.1 Mean Photon Number
Using the BS transformations, the mean values for $|n\rangle_{a}|\alpha\rangle_{b}$ is:
$\langle\mathcal{N}_{a}\rangle=\langle a_{a}^{\dagger}a_{a}\rangle=n,\quad\langle\mathcal{N}_{b}\rangle=\langle a_{b}^{\dagger}a_{b}\rangle=|\alpha|^{2}$ (3.5)
Thus:
$\langle\mathcal{N}_{c}\rangle$ $=n|t|^{2}+|r|^{2}|\alpha|^{2},\quad\langle\mathcal{N}_{d}\rangle=n|r|^{2}+|t|^{2}|\alpha|^{2}.$ (3.6)
Where the total photon number is conserved:
$\langle\mathcal{N}_{c}\rangle+\langle\mathcal{N}_{d}\rangle=n+|\alpha|^{2}.$ (3.7)
3.2 Photon Number Variance
For the variance in mode $c$ $(\Delta\mathcal{N}_{c})^{2}=\langle\mathcal{N}_{c}^{2}\rangle-\langle\mathcal{N}_{c}\rangle^{2}$ , we first calculate $\langle\mathcal{N}_{c}^{2}\rangle$
$\langle\mathcal{N}^{2}_{c}\rangle$ $=\Big{\langle}\Big{(}|t|^{2}n+|r|^{2}|\alpha|^{2}+itr^{*}a^{\dagger}_{a}a_{b}-it^{*}ra^{\dagger}_{b}a_{a}\Big{)}^{2}\Big{\rangle}$
$=|t|^{4}n^{2}+2|t|^{2}|r|^{2}n|\alpha|^{2}+\langle|r|^{4}\underbrace{(a^{\dagger}_{b}a_{b})^{2}}_{|\alpha|^{4}+|\alpha|^{2}}\rangle$
$\quad+\Big{\langle}|t|^{2}|r|^{2}\Big{(}a^{\dagger}_{a}\underbrace{a_{b}a^{\dagger}_{b}}_{\mathcal{N}_{b}+1}a_{a}+a^{\dagger}_{b}\underbrace{a_{a}a^{\dagger}_{a}}_{\mathcal{N}_{a}+1}a_{b}\Big{)}\Big{\rangle}$
$=\langle\mathcal{N}_{c}\rangle^{2}++|r|^{4}|\alpha|^{2}+|t|^{2}|r|^{2}(2n|\alpha|^{2}+n+|\alpha|^{2}),$ (3.8)
hence
$(\Delta\mathcal{N}_{c})^{2}=|r|^{4}|\alpha|^{2}+|t|^{2}|r|^{2}(2n|\alpha|^{2}+n+|\alpha|^{2}).$ (3.9)
3.3 Mandel $Q$ Parameter
For mode $c$:
$Q_{c}$ $=\frac{|r|^{4}|\alpha|^{4}+|t|^{2}|r|^{2}(n+|\alpha|^{2}+2n|\alpha|^{2})}{n|t|^{2}+|r|^{2}|\alpha|^{2}}-1,$ (3.10)
This parameter can be negative (sub-Poissonian) when the Fock state dominates, or positive (super-Poissonian) when the coherent state dominates, reflecting the hybrid quantum-classical nature of the output.
Figure 4: Mandel $Q_{c}$ parameter for mode $c$ of a 50:50 BS with Fock-coherent state input of the type $|n\rangle_{a}|\alpha\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$. (a) For $n=0$ (vacuum in mode $a$), $Q_{c}=0$ (Poissonian statistics) as the output remains coherent; (b) For $n\geq 1$, $Q_{c}$ exhibits sub-Poissonian statistics ($Q_{c}<0$) at small $\alpha$ due to Fock state antibunching, with $Q_{c}\rightarrow-1/2$ when $\alpha\rightarrow 0$, reflecting single-mode Fock state variance suppression; (c) As $\alpha$ increases, $Q_{c}>0$ (super-Poissonian) due to coherent state phase randomness. Numerical parameters: $\alpha\in[0,3]$, $n\in\{0,1,2,5,10\}$.
Figure 5: $g^{(2)}(0)$ for mode $c$ of a 50:50 BS with Fock-Fock input of the type $|n\rangle_{a}|m\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$. As $|\alpha|\to 0$, the behavior approaches that of a single Fock state input, while for large $|\alpha|$, the statistics become increasingly classical.
3.4 Second-order coherence function $g^{2}(0)$
For the $c$-mode, the $g$ parameter is:
$g^{2}_{c}(0)$ $=1+\frac{1}{n|t|^{2}+|r|^{2}|\alpha|^{2}}$
$\quad\times\Bigg{(}\frac{|r|^{4}|\alpha|^{4}+|t|^{2}|r|^{2}(n+|\alpha|^{2}+2n|\alpha|^{2})}{n|t|^{2}+|r|^{2}|\alpha|^{2}}-1\Bigg{)}.$ (3.11)
$Q$ and $g$ parameters are plotted in figures (4) and (5), respectively.
3.5 Special Cases
3.5.1 Vacuum and coherent state $|0\rangle_{a}|\alpha\rangle_{b}$.
Output state:
$|out\rangle_{cd}=|ir\alpha\rangle_{c}|t\alpha\rangle_{d}.$ (3.12)
Statistics remain Poissonian:
$\langle\mathcal{N}_{c}\rangle=(\Delta\mathcal{N}_{c})^{2}=|r|^{2}|\alpha|^{2},\ Q_{c}=0,\ g^{2}_{c}(0)=1.$ (3.13)
3.5.2 Single-photon and coherent state $|1\rangle_{a}|\alpha\rangle_{b}$.
Output state:
$|out\rangle_{cd}=ta_{c}^{\dagger}|ir\alpha\rangle_{c}|t\alpha\rangle_{d}+ira_{d}^{\dagger}|ir\alpha\rangle_{c}|t\alpha\rangle_{d}.$ (3.14)
With:
$\langle\mathcal{N}_{c}\rangle$ $=|t|^{2}+|r|^{2}|\alpha|^{2},$
$(\Delta\mathcal{N}_{c})^{2}$ $=|r|^{4}|\alpha|^{2}+|t|^{2}|r|^{2}(1+|\alpha|^{2}+2|\alpha|^{2}),$
$Q_{c}$ $=\frac{|\alpha|^{4}+3|\alpha|^{2}+1}{2(1+|\alpha|^{2})}-1,$
$g^{2}_{c}(0)$ $=1+\frac{|\alpha|^{4}+|\alpha|^{2}-1}{(1+|\alpha|^{2})^{2}}.$
4 Coherent State Inputs $|\alpha\rangle_{a}|\beta\rangle_{b}$
We consider the case where both input modes are in coherent states:
$|in\rangle_{ab}=|\alpha\rangle_{a}|\beta\rangle_{b}=\mathcal{D}_{a}(\alpha)\mathcal{D}_{b}(\beta)|0\rangle_{a}|0\rangle_{b},$ (4.1)
where $\mathcal{D}(\xi)=\exp(\xi a^{\dagger}-\xi^{*}a)$ is the displacement operator for complex amplitude $\xi$.
The BS transforms coherent states into coherent states. Using the transformation property:
$U_{BS}\mathcal{D}_{a}(\alpha)\mathcal{D}_{b}(\beta)U_{BS}^{\dagger}=\mathcal{D}_{c}(t\alpha+ir\beta)\mathcal{D}_{d}(ir\alpha+t\beta),$ (4.2)
the output state is:
$|out\rangle_{cd}=|t\alpha+ir\beta\rangle_{c}|ir\alpha+t\beta\rangle_{d}.$ (4.3)
This shows the BS linearly transforms the input amplitudes $(\alpha,\beta)$ to output amplitudes $(t\alpha+ir\beta,ir\alpha+t\beta)$.
4.1 Mean Photon Number
Since coherent states remain coherent under BS transformation:
$\langle\mathcal{N}_{c}\rangle=|t\alpha+ir\beta|^{2},\quad\langle\mathcal{N}_{d}\rangle=|ir\alpha+t\beta|^{2}.$ (4.4)
Where the total photon number is conserved:
$\langle\mathcal{N}_{c}\rangle+\langle\mathcal{N}_{d}\rangle=|\alpha|^{2}+|\beta|^{2}.$ (4.5)
4.2 Photon Number Variance
For coherent states, photon number variance equals the mean (since it is Poissonian) [1, 2]:
$(\Delta\mathcal{N}_{c})^{2}=\langle\mathcal{N}_{c}\rangle,\quad(\Delta\mathcal{N}_{d})^{2}=\langle\mathcal{N}_{d}\rangle.$ (4.6)
4.3 Mandel Q parameter
The Mandel $Q$ parameter is zero for both modes:
$Q_{c}=Q_{d}=0,$ (4.7)
confirming Poissonian statistics are preserved.
4.4 Second-order coherence function $g^{2}(0)$
Since coherent states follow Poisson statistics, the second-order coherent parameter is 1 for both modes.
4.5 Special Cases
4.5.1 One port in vacuum ($\beta=0$) $|\alpha\rangle_{a}|0\rangle_{b}$
We recover equation (3.12) exchanging $t$ for $r.$
4.5.2 Balanced inputs and BS ($\alpha=\beta$, $|t|^{2}=|r|^{2}=1/2$).
Output amplitudes:
$t\alpha+ir\beta$ $=\frac{1}{\sqrt{2}}(\alpha+i\alpha)=\frac{(1+i)\alpha}{\sqrt{2}},$
$ir\alpha+t\beta$ $=\frac{1}{\sqrt{2}}(i\alpha+\alpha)=\frac{(1+i)\alpha}{\sqrt{2}}.$ (4.8)
Thus:
$|out\rangle_{cd}=\left|\frac{(1+i)\alpha}{\sqrt{2}}\right\rangle_{c}\left|\frac{(1+i)\alpha}{\sqrt{2}}\right\rangle_{d}.$ (4.9)
Each output mode has mean photon number $|\alpha|^{2}$.
5 Conclusions
In this work, we have presented a theoretical analysis of the photon statistics resulting from the interference of quantum light fields at an ideal lossless beamsplitter (BS). By exploring a broad range of input states, namely Fock–Fock, Fock–coherent, and coherent–coherent configurations, we derived explicit analytical expressions for the output states and computed fundamental statistical quantities, including the mean photon number, photon number variance, Mandel $Q$ parameter, and $g$ function. These calculations were carried out using standard operator techniques and the well-known properties of coherent and Fock states, which highlight the quantum structure underlying the field transformations induced by the BS.
Our results show the close interplay between interference and quantum statistics in shaping the output photon distributions. In particular, for specific Fock state inputs, we recovered nonclassical features such as sub-Poissonian statistics (e.g., when sending vacuum and single-photon states in the input modes) and the Hong–Ou–Mandel interference effect (for a single-photon ejected at each input mode of a balanced BS), which originates from the indistinguishability and bosonic nature of photons. For hybrid inputs combining Fock and coherent states, we analyzed how the coherent component partially preserves its classical-like statistical behavior while still exhibiting interference effects due to the quantum nature of the Fock component. In the case of purely coherent inputs, we confirmed that the output states remain coherent and retain their Poissonian photon number distributions, with the interference being fully determined by the adjustable BS coefficients $t$ and $r$.
In addition to these statistical moments, we also computed the normalized second-order correlation function at zero delay, $g^{(2)}(0)$, for each output mode. This quantity provides a deeper insight into the photon bunching and antibunching behavior of the output fields, revealing regimes of classical and nonclassical light. For Fock–Fock inputs, we recovered exact expressions demonstrating antibunching effects for single-photon inputs and bunching for higher photon numbers. In the case of Fock–coherent interference, our results captured how the presence of a coherent component smooths out the nonclassical correlations, interpolating between sub-Poissonian and Poissonian statistics depending on the relative intensities. These analytical expressions for $g^{(2)}(0)$ help elucidate the transition from quantum to classical behavior in beamsplitter outputs, and offer practical utility in characterizing sources and designing experiments in quantum optics.
Beyond these specific cases, our study provides a unified perspective on how different statistical regimes, ranging from purely quantum (sub-Poissonian) to classical-like (Poissonian), emerge from the linear mixing of quantum states at a BS. This viewpoint is particularly relevant for understanding fundamental experiments in quantum optics, such as single-photon interference and coherent state splitting, as well as for interpreting modern applications in quantum information processing and photonic technologies.
Finally, by presenting the derivations and results in a clear and accessible way, we aim to provide a complementary reference and an intuitive framework for students and researchers interested in the quantum statistical properties of light in a beamsplitter. We hope that this study can serve as a starting point for more advanced investigations of quantum interference phenomena and their applications.
Acknowledgements.
J. A. T. S. is grateful for the encouragement of Prof. Danieverton Moretti and the financial support provided by UFCG.